OpenStudy (anonymous):

Given the equation x^2+y^2-8x-6y=0 (a) Write the equation in standard form. (b) State the center, radius, and intercepts. (c) Submit the graph. Please show all of your work.

OpenStudy (anonymous):

x^2+y^2-8x-6y=0 this is the equation of circle, rearranging them, x^2-8x+y^2-6y=0 we complete the squares x^2-8x +y^2-6y =0 x^2-8x +(8/2)^2 +y^2-6y+(6/2)^2 =0 +(8/2)^2 +(6/2)^2 x^2-8x +16 +y^2-6y+9 =0 +16 +9 x^2-8x+16 +y^2-6y+9 =25 (x-4)^2+(y-3)^2=25 ans....equation of circle

OpenStudy (anonymous):

from the formula of the eq of circle (x-h)^2+(y-k)^2=r^2 (x-4)^2+(y-3)^2=5^2 we see that h=4,k=3 and r=5 ans.........

OpenStudy (anonymous):

centre (h,k)=(4,3) with radius r=5 ans

OpenStudy (anonymous):

good luck now sclowers

OpenStudy (anonymous):

thank you mark!

OpenStudy (anonymous):

yw, have fun